OpenStudy (anonymous):

find the limit (if it exists). lim (1-[I -x/2l]) x->1

OpenStudy (lgbasallote):

\[\huge \lim_{x\rightarrow 1} (1 - [1 - \frac{x}{21}] )\] is that your question?

OpenStudy (anonymous):

no i was trying to show the greatest integer symbol/sign

OpenStudy (lgbasallote):

oh..that

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

as \(x\) gets close to one, \(\frac{x}{2}\) gets close to \(\frac{1}{2}\) so greatest integer(\frac{x}{2} is 0 in an interval about 1

OpenStudy (anonymous):

i meant greatest integer of \(\frac{x}{2}\) is 0 in an interval about 1